I'm trying to find the table of ordered pairs.
I can't solve, and need help please.
Equation: X+3y=18
X y
0 6
0
8

Answer by Edwin McCravy(20055)

X+3y=18 

X y
0 6
0
8

Substitute 0 for X in the equation X+3y=18

X+3y=18
0+3y=18
  3y=18
Divide both sides by 3
   y=6

You already have a 6 there.

X y
0 6
0
8

I don't know why you have a 0 twice.  
That's also 6 for y, so we'll put 6 there too

X y
0 6
0 6 
8

The 8 is the only one left.
Substitute 8 for X in

X+3y=18
8+3y=18
Subtract 8 from both sides
  3y=10
Divide both sides by 3
   y=10/3 = 3 1/3

So you can put either 10/3 or 3 1/3

X y
0 6
0 6 
8 10/3 or 3 1/3

To graph those two ordered pairs, (0,6), and (8,3 1/3).
To get 3 1/3, you have to guess that's its even with the
point on the y-axis that's about 1/3 of the way between
3 and 4.  Notice the green line that I've drawn going
over to the y-axis to show you that it's about 1/3 of
the way between 3 and 4. I've also drawn a green line
down to the x-axis to show you that the point (8, 3 1/3)
is directly above it's x-coordinate value 8 on the x-axis,
and directly beside is y-coordinate 3 1/3 on the y-axis.



Now I'll erase the green lines, because they aren't part of the graph.
Now we get a ruler and draw a line through those two points, and
that line is the graph of the equation X+3y=18.

 

Edwin